# Widths of embeddings of Gaussian Sobolev spaces

### 14 Apr 2026 v1

@@ 2 Gaussian and fractional Gaussian Sobolev spaces
Before introducing fractional Gaussian Sobolev spaces on $\mathbb{R}^{d}$, we define the corresponding fractional Sobolev spaces on $\mathbb{U}^{d}$, where $\mathbb{U}=[a,b]$ or $\mathbb{U}=\mathbb{R}$. As usual, $L_{p}(\mathbb{U}^{d})$ denotes the Lebesgue space of functions on $\mathbb{U}^{d}$ equipped with the standard $L_{p}$-norm. For $s\in\mathbb{N}$, the space $W^{s}_{p}(\mathbb{U}^{d})$ is defined analogously to $W^{s}_{p}(\mathbb{R}^{d},\gamma)$ by replacing $L_{p}(\mathbb{R}^{d},\gamma)$ in (2.1) with $L_{p}(\mathbb{U}^{d})$.
@@
Ornstein–Uhlenbeck operator in ${\mathbb{R}}^{d}$ is a second-order differential operator given by
@@
In the following we collect some properties of the kernel $K_{\sigma}({\boldsymbol{x}},{\boldsymbol{y}})$ which is useful to formulate our result.
@@
For ${\boldsymbol{x}},{\boldsymbol{y}}\in{\mathbb{R}}^{d}$ and $\sigma\in(0,1)$ we have
@@
This lemma was proved in [2, Lemma 2.8]. We have some further estimates.
@@
    If $|{\boldsymbol{x}}-{\boldsymbol{y}}|>1$, then it holds
    $K_{\sigma}({\boldsymbol{x}},{\boldsymbol{y}})\leq C\exp\bigg(\frac{|{\boldsymbol{x}}|^{2}+|{\boldsymbol{y}}|^{2}}{4}\bigg).$
@@
    If $|{\boldsymbol{x}}-{\boldsymbol{y}}|\leq 1$ and $t_{0}>0$, then
    $K_{\sigma}({\boldsymbol{x}},{\boldsymbol{y}})\geq C_{t_{0}}\exp\bigg(\frac{|{\boldsymbol{x}}|^{2}+|{\boldsymbol{y}}|^{2}}{4}\Big(1-\frac{1-e^{-t_{0}}}{1+e^{-t_{0}}}\Big)\bigg)\frac{1}{|{\boldsymbol{x}}-{\boldsymbol{y}}|^{d+\sigma}},$
    where $C_{t_{0}}$ is a positive constant depending only on $t_{0}$.
@@
$\leq C_{1}\exp\bigg(\frac{|{\boldsymbol{x}}|^{2}+|{\boldsymbol{y}}|^{2}}{4}\bigg)\bigg(\int_{1}^{\infty}\frac{{\rm d}t}{t^{\frac{\sigma}{2}+1}},+\int_{0}^{1}\frac{\exp\big(-\frac{|{\boldsymbol{x}}-{\boldsymbol{y}}|^{2}}{4t}\big)}{(2t)^{d/2}t^{\frac{\sigma}{2}+1}}{\rm d}t\bigg)$
@@
Performing a change of variables in the second term on the right-hand side we get
@@
Performing a change of variables in the second term on the right-hand side and using $|{\boldsymbol{x}}-{\boldsymbol{y}}|\leq 1$ we get
@@
$\geq\frac{2^{\sigma+d}}{|{\boldsymbol{x}}-{\boldsymbol{y}}|^{d+\sigma}}\int_{\frac{1}{4t_{0}}}^{\infty}e^{-\xi}\xi^{\frac{\sigma}{2}+\frac{d}{2}-1}{\rm d}\xi.$
@@
$\|f\|_{W_{p,G}^{s}({\mathbb{R}}^{d},\gamma)}\leq C\|f\|_{W_{p}^{s}({\mathbb{R}}^{d},\gamma)}.$
@@
$\|f\|_{H^{s}(\mathbb{R}^{d},\gamma)}:=\Bigg(\sum_{{\boldsymbol{k}}\in\mathbb{N}_{0}^{d}}(1+|{\boldsymbol{k}}|_{1})^{s}|\widehat{f}({\boldsymbol{k}})|^{2}\Bigg)^{1/2}<\infty$
@@
Hence we have for ${\boldsymbol{\alpha}}\in{\mathbb{N}}_{0}^{d}$ with $|{\boldsymbol{\alpha}}|_{1}\leq s$
$\int_{{\mathbb{R}}^{d}}|D^{\boldsymbol{\alpha}}f({\boldsymbol{x}})|^{2}{\rm d}\gamma({\boldsymbol{x}})=\sum_{{\boldsymbol{k}}\geq{\boldsymbol{\alpha}}}\frac{{\boldsymbol{k}}!}{\left({\boldsymbol{k}}-{\boldsymbol{\alpha}}\right)!}|\widehat{f}({\boldsymbol{k}})|^{2}\asymp\sum_{{\boldsymbol{k}}\geq{\boldsymbol{\alpha}}}\big(k_{1}^{\alpha_{1}}\cdots k_{d}^{\alpha_{d}}\big)|\widehat{f}({\boldsymbol{k}})|^{2},$
with the convention $0^{0}=1$. This derives
$\|f\|_{W^{s}_{p}(\mathbb{R}^{d},\gamma)}\asymp\Bigg(\sum_{{\boldsymbol{k}}\in\mathbb{N}_{0}^{d}}(1+|{\boldsymbol{k}}|_{1})^{s}|\widehat{f}({\boldsymbol{k}})|^{2}\Bigg)^{1/2}$
which proves the case $s\in{\mathbb{N}}$. Now we consider the case $s\not\in{\mathbb{N}}$. We us the following formula for $v\in W^{\sigma}_{2}({\mathbb{R}}^{d},\gamma)$
$\int_{\mathbb{R}^{d}}{\rm d}\gamma({\boldsymbol{x}})\int_{\mathbb{R}^{d}}|v({\boldsymbol{x}})-v({\boldsymbol{y}})|^{2}\,K_{2\sigma}({\boldsymbol{x}},{\boldsymbol{y}})\,{\rm d}\gamma({\boldsymbol{y}})$ $=2\Gamma(-\sigma)\int_{\mathbb{R}}v({\boldsymbol{x}})\,(-\Delta_{\gamma})^{\sigma}v({\boldsymbol{x}})\,{\rm d}\gamma({\boldsymbol{x}})$ (2.8)
which can be found in [2, page 6]. For $f\in W^{s}_{p}({\mathbb{R}}^{d},\gamma)$ we put $v=D^{\boldsymbol{\alpha}}f$ with $|{\boldsymbol{\alpha}}|_{1}=\bar{s}$. By using (2.8) we get
${}_{W_{2}^{s}(\mathbb{R}^{d},\gamma)}^{2}$ $=2\Gamma(-\tilde{s})\int_{\mathbb{R}}v({\boldsymbol{x}})\,(-\Delta_{\gamma})^{\tilde{s}}v({\boldsymbol{x}})\,{\rm d}\gamma({\boldsymbol{x}})$ (2.9)
$=2\Gamma(-\tilde{s})\sum_{{\boldsymbol{k}}\in{\mathbb{N}}_{0}^{d}}\hat{v}({\boldsymbol{k}})\int_{\mathbb{R}}v({\boldsymbol{x}})\,(-\Delta_{\gamma})^{\tilde{s}}H_{\boldsymbol{k}}({\boldsymbol{x}})\,{\rm d}\gamma({\boldsymbol{x}})$
$=2\Gamma(-\tilde{s})\sum_{{\boldsymbol{k}}\in{\mathbb{N}}_{0}^{d}}\hat{v}({\boldsymbol{k}})\int_{\mathbb{R}}v({\boldsymbol{x}})\,(k_{1}^{\tilde{s}}+\ldots+k_{d}^{\tilde{s}})H_{\boldsymbol{k}}({\boldsymbol{x}})\,{\rm d}\gamma({\boldsymbol{x}})$
$=2\Gamma(-\tilde{s})\sum_{{\boldsymbol{k}}\in{\mathbb{N}}_{0}^{d}}(k_{1}^{\tilde{s}}+\ldots+k_{d}^{\tilde{s}})|\hat{v}({\boldsymbol{k}})|^{2}.$
@@
$\hat{v}({\boldsymbol{k}})$ $=\int_{{\mathbb{R}}^{d}}D^{\boldsymbol{\alpha}}f({\boldsymbol{x}})H_{\boldsymbol{k}}({\boldsymbol{x}}){\rm d}\gamma({\boldsymbol{x}})$
@@
$f]_{W_{2}^{s}(\mathbb{R}^{d},\gamma)}^{2}=2\Gamma(-\tilde{s})\sum_{{\boldsymbol{k}}\in{\mathbb{N}}_{0}^{d}}(k_{1}^{\tilde{s}}+\ldots+k_{d}^{\tilde{s}})\frac{({\boldsymbol{\alpha}}+{\boldsymbol{k}})!}{{\boldsymbol{k}}!}|\widehat{f}({\boldsymbol{\alpha}}+{\boldsymbol{k}})|^{2}.$
@@ 3 Approximation for Gaussian Sobolev spaces
Let $\Omega\subset\mathbb{R}^{d}$ be a domain, and let $Y$ be a Banach space of functions defined on $\Omega$. Let $F\subset Y$ be a compact set and let $n\in\mathbb{N}$. Given sampling points $\{{\boldsymbol{x}}_{i}\}_{i=1}^{n}\subset\Omega$, we aim to approximately recover a function $f\in F$ from the sampled values ${f({\boldsymbol{x}}_{i})}_{i=1}^{n}$ by means of a linear sampling algorithm defined by
@@
where $\{\varphi_{i}\}_{i=1}^{n}$ is a family of functions in $Y$. For the sake of generality, we permit repetitions both in the set of sampling points $\{{\boldsymbol{x}}_{i}\}_{i=1}^{n}$ and in the collection of functions $\{\varphi_{i}\}_{i=1}^{n}$. For each $n\in\mathbb{N}$, the sampling $n$-width of the set $F$ in $Y$ is defined as follows:
@@
If $T$ is a linear bounded operator from two Banach spaces $X$ and $Y$. Then
$d_{n}(T)=d_{n}(T(B_{X}),Y),\ \ a_{n}(T)=\lambda_{n}(T(B_{X}),Y).$
Here, $B_{X}$ denotes the closed unit ball in $X$ and $d_{n}(T)$, $a_{n}(T)$ are Kolmogorov and approximation numbers of $T$, respectively. For a detailed exposition of the theory of widths and related $s$-numbers, we refer the reader to the monographs [15, 13, 14].
Denote by $\tilde{L}_{q}({\mathbb{U}}^{d})$ and $\tilde{W}^{\alpha}_{p}({\mathbb{U}}^{d})$ the subspaces of $L_{q}({\mathbb{U}}^{d})$ and $W^{s}_{p}({\mathbb{U}}^{d})$, respectively, of all functions $f$ which can be extended to the whole ${\mathbb{R}}^{d}$ as periodic functions in each variable. The following proposition is crucial for the formulation of our main results.
@@
Choose $q_{*}$ and $p_{*}$ such that $1\leq q<q_{*}<p_{*}<p$. When ${\boldsymbol{x}}\in{\mathbb{I}}^{d}_{\theta,{\boldsymbol{k}}}$ we have $e^{\frac{|{\boldsymbol{x}}|^{2}}{2}}\leq e^{\frac{p|{\boldsymbol{k}}|^{2}}{2p_{*}}}$. Therefore,
$\|f_{\boldsymbol{k}}(\cdot+{\boldsymbol{k}})\|_{W_{p}^{\bar{s}}({\mathbb{I}}^{d}_{\theta})}\leq(2\pi)^{\frac{d}{2p}}e^{\frac{|{\boldsymbol{k}}|^{2}}{2p_{*}}}\|f\|_{W^{s}_{p}({\mathbb{R}}^{d},\gamma)}.$
@@
$\int_{{\mathbb{I}}^{d}_{\theta}}\int_{{\mathbb{I}}^{d}_{\theta}}\frac{\big|D^{\boldsymbol{\alpha}}f_{\boldsymbol{k}}({\boldsymbol{x}}+{\boldsymbol{k}})-D^{\boldsymbol{\alpha}}f({\boldsymbol{y}}+{\boldsymbol{k}})\big|^{p}}{|{\boldsymbol{x}}-{\boldsymbol{y}}|^{d+\tilde{s}p}}{\rm d}{\boldsymbol{x}}{\rm d}{\boldsymbol{y}}$
$\leq C_{t_{0}}\int_{{\mathbb{I}}^{d}_{\theta,{\boldsymbol{k}}}}\int_{{\mathbb{I}}^{d}_{\theta,{\boldsymbol{k}}}}\big|D^{\boldsymbol{\alpha}}f_{\boldsymbol{k}}({\boldsymbol{x}})-D^{\boldsymbol{\alpha}}f({\boldsymbol{y}})\big|^{p}e^{-\frac{|{\boldsymbol{x}}|^{2}+|{\boldsymbol{y}}|^{2}}{4}\big(1-\frac{1-e^{-t_{0}}}{1+e^{-t_{0}}}\big)}K_{p\tilde{s}}({\boldsymbol{x}},{\boldsymbol{y}}){\rm d}{\boldsymbol{x}}{\rm d}{\boldsymbol{y}}$
$\ll C_{t_{0}}\int_{{\mathbb{I}}^{d}_{\theta,{\boldsymbol{k}}}}\int_{{\mathbb{I}}^{d}_{\theta,{\boldsymbol{k}}}}\big|D^{\boldsymbol{\alpha}}f_{\boldsymbol{k}}({\boldsymbol{x}})-D^{\boldsymbol{\alpha}}f({\boldsymbol{y}})\big|^{p}e^{\frac{|{\boldsymbol{x}}|^{2}+|{\boldsymbol{y}}|^{2}}{4}\big(1+\frac{1-e^{-t_{0}}}{1+e^{-t_{0}}}\big)}K_{p\tilde{s}}({\boldsymbol{x}},{\boldsymbol{y}}){\rm d}\gamma({\boldsymbol{x}}){\rm d}\gamma({\boldsymbol{y}})$
$\ll C_{t_{0}}e^{\frac{p|{\boldsymbol{k}}|^{2}}{2p_{*}}\big(1+\frac{1-e^{-t_{0}}}{1+e^{-t_{0}}}\big)}\|f\|_{W^{s}_{p}({\mathbb{R}}^{d},\gamma)}.$
Therefore, we find that
$\|f_{\boldsymbol{k}}(\cdot+{\boldsymbol{k}})\|_{W^{s}_{p}({\mathbb{I}}^{d}_{\theta})}\ll C_{t_{0}}e^{\frac{|{\boldsymbol{k}}|^{2}}{2p_{*}}\big(1+\frac{1-e^{-t_{0}}}{1+e^{-t_{0}}}\big)}\|f\|_{W^{s}_{p}({\mathbb{R}}^{d},\gamma)}.$
@@
$\|f_{{\boldsymbol{k}}}(\cdot+{\boldsymbol{k}})\varphi_{{\boldsymbol{k}}}(\cdot+{\boldsymbol{k}})\|_{W^{s}_{p}({\mathbb{I}}^{d}_{\theta})}$ $\ll C_{t_{0}}\|f_{{\boldsymbol{k}}}(\cdot+{\boldsymbol{k}})\|_{W^{s}_{p}({\mathbb{I}}^{d}_{\theta})}\left\|{\varphi_{\boldsymbol{k}}(\cdot+{\boldsymbol{k}})}\right\|_{C^{\kappa}({\mathbb{R}}^{d})}$
$\ll C_{t_{0}}e^{\frac{|{\boldsymbol{k}}|^{2}}{2p_{*}}\big(1+\frac{1-e^{-t_{0}}}{1+e^{-t_{0}}}\big)}\|f\|_{W^{s}_{p}({\mathbb{R}}^{d},\gamma)}.$

### 18 Apr 2026 v2

@@ 2 Gaussian and fractional Gaussian Sobolev spaces
Before introducing fractional Gaussian Sobolev spaces on $\mathbb{R}^{d}$, we define the classical fractional Sobolev spaces on $\mathbb{U}^{d}$, where $\mathbb{U}=[a,b]$ or $\mathbb{U}=\mathbb{R}$. As usual, $L_{p}(\mathbb{U}^{d})$ denotes the Lebesgue space of functions on $\mathbb{U}^{d}$ equipped with the standard $L_{p}$-norm. For $s\in\mathbb{N}$, the space $W^{s}_{p}(\mathbb{U}^{d})$ is defined analogously to $W^{s}_{p}(\mathbb{R}^{d},\gamma)$ by replacing $L_{p}(\mathbb{R}^{d},\gamma)$ in (2.1) with $L_{p}(\mathbb{U}^{d})$.
@@
Ornstein–Uhlenbeck operator in ${\mathbb{R}}^{d}$ is a second-order partial differential operator given by
@@
In the following we collect some properties of the kernel $K_{\sigma}({\boldsymbol{x}},{\boldsymbol{y}})$ which is useful to formulate our results.
@@
For ${\boldsymbol{x}},{\boldsymbol{y}}\in{\mathbb{R}}^{d}$ and $\sigma>0$ we have
@@
This lemma was proved in [2, Lemma 2.8]. We have some further estimates of the kernel $K_{\sigma}({\boldsymbol{x}},{\boldsymbol{y}})$.
@@
    If $|{\boldsymbol{x}}-{\boldsymbol{y}}|\geq M$, then it holds
    $K_{\sigma}({\boldsymbol{x}},{\boldsymbol{y}})\leq C_{M}\exp\bigg(\frac{|{\boldsymbol{x}}|^{2}+|{\boldsymbol{y}}|^{2}}{4}\bigg).$
@@
    Let ${\boldsymbol{\ell}}\in{\mathbb{R}}^{d}$ and $t_{0}>0$. If ${\boldsymbol{x}},{\boldsymbol{y}}\in[0,1]^{d}+{\boldsymbol{\ell}}$, then
    $K_{\sigma}({\boldsymbol{x}},{\boldsymbol{y}})\geq C_{d,t_{0}}\exp\bigg(\frac{|{\boldsymbol{x}}|^{2}+|{\boldsymbol{y}}|^{2}}{4}\Big(1-\frac{1-e^{-t_{0}}}{1+e^{-t_{0}}}\Big)\bigg)\frac{1}{|{\boldsymbol{x}}-{\boldsymbol{y}}|^{d+\sigma}},$
    where $C_{d,t_{0}}$ is a positive constant depending only on $d$ and $t_{0}$.
@@
$\leq C_{1}\exp\bigg(\frac{|{\boldsymbol{x}}|^{2}+|{\boldsymbol{y}}|^{2}}{4}\bigg)\bigg(\int_{1}^{\infty}\frac{{\rm d}t}{t^{\frac{\sigma}{2}+1}}+\int_{0}^{1}\frac{\exp\big(-\frac{|{\boldsymbol{x}}-{\boldsymbol{y}}|^{2}}{4t}\big)}{(2t)^{d/2}t^{\frac{\sigma}{2}+1}}{\rm d}t\bigg)$
@@
Performing change of variable in the second term on the right-hand side we get
@@
Performing a change of variables in the second term on the right-hand side and using $|{\boldsymbol{x}}-{\boldsymbol{y}}|\leq\sqrt{d}$ we get
@@
$\geq\frac{2^{\sigma+d}}{|{\boldsymbol{x}}-{\boldsymbol{y}}|^{d+\sigma}}\int_{\frac{\sqrt{d}}{4t_{0}}}^{\infty}e^{-\xi}\xi^{\frac{\sigma}{2}+\frac{d}{2}-1}{\rm d}\xi.$
@@
$\|f\|_{W_{p,G}^{s}({\mathbb{R}}^{d},\gamma)}\leq C\|f\|_{W_{p}^{s}({\mathbb{R}}^{d},\gamma)}$
and hence
$W_{p}^{s}(\mathbb{R}^{d},\gamma)\hookrightarrow W^{s}_{p,G}({\mathbb{R}}^{d},\gamma).$
@@
$\|f\|_{H^{s}(\mathbb{R}^{d},\gamma)}:=\Bigg(\sum_{{\boldsymbol{k}}\in\mathbb{N}_{0}^{d}}(1+|{\boldsymbol{k}}|_{1})^{s}|\widehat{f}({\boldsymbol{k}})|^{2}\Bigg)^{1/2}<\infty.$
@@
Hence, we get for ${\boldsymbol{\alpha}}\in{\mathbb{N}}_{0}^{d}$ with $|{\boldsymbol{\alpha}}|_{1}\leq s$
$\int_{{\mathbb{R}}^{d}}|D^{\boldsymbol{\alpha}}f({\boldsymbol{x}})|^{2}{\rm d}\gamma({\boldsymbol{x}})=\sum_{{\boldsymbol{k}}\geq{\boldsymbol{\alpha}}}\frac{{\boldsymbol{k}}!}{\left({\boldsymbol{k}}-{\boldsymbol{\alpha}}\right)!}|\widehat{f}({\boldsymbol{k}})|^{2}\asymp\sum_{{\boldsymbol{k}}\geq{\boldsymbol{\alpha}}}\big(k_{1}^{\alpha_{1}}\cdot\ldots\cdot k_{d}^{\alpha_{d}}\big)|\widehat{f}({\boldsymbol{k}})|^{2},$
with the convention $0^{0}=1$. This derives the case $s\in{\mathbb{N}}$
$\|f\|_{W^{s}_{p}(\mathbb{R}^{d},\gamma)}\asymp\Bigg(\sum_{{\boldsymbol{k}}\in\mathbb{N}_{0}^{d}}(1+|{\boldsymbol{k}}|_{1})^{s}|\widehat{f}({\boldsymbol{k}})|^{2}\Bigg)^{1/2}.$
Now we consider the case $s\not\in{\mathbb{N}}$. We use the following formula for $v\in W^{\sigma}_{2}({\mathbb{R}}^{d},\gamma)$
$\int_{\mathbb{R}^{d}}{\rm d}\gamma({\boldsymbol{x}})\int_{\mathbb{R}^{d}}|v({\boldsymbol{x}})-v({\boldsymbol{y}})|^{2}\,K_{2\sigma}({\boldsymbol{x}},{\boldsymbol{y}})\,{\rm d}\gamma({\boldsymbol{y}})$ $=2\Gamma(-\sigma)\int_{\mathbb{R}^{d}}v({\boldsymbol{x}})\,(-\Delta_{\gamma})^{\sigma}v({\boldsymbol{x}})\,{\rm d}\gamma({\boldsymbol{x}})$ (2.8)
which can be found in [2, page 6]. For $f\in W^{s}_{p}({\mathbb{R}}^{d},\gamma)$ we put $v=D^{\boldsymbol{\alpha}}f$ with ${\boldsymbol{\alpha}}\in{\mathbb{N}}_{0}^{d}$ and $|{\boldsymbol{\alpha}}|_{1}=\bar{s}$. By using (2.8) we get
${}_{W_{2}^{s}(\mathbb{R}^{d},\gamma)}^{2}$ $=2\Gamma(-\tilde{s})\int_{\mathbb{R}^{d}}v({\boldsymbol{x}})\,(-\Delta_{\gamma})^{\tilde{s}}v({\boldsymbol{x}})\,{\rm d}\gamma({\boldsymbol{x}})$ (2.9)
$=2\Gamma(-\tilde{s})\sum_{{\boldsymbol{k}}\in{\mathbb{N}}_{0}^{d}}\widehat{v}({\boldsymbol{k}})\int_{\mathbb{R}^{d}}v({\boldsymbol{x}})\,(-\Delta_{\gamma})^{\tilde{s}}H_{\boldsymbol{k}}({\boldsymbol{x}})\,{\rm d}\gamma({\boldsymbol{x}})$
$=2\Gamma(-\tilde{s})\sum_{{\boldsymbol{k}}\in{\mathbb{N}}_{0}^{d}}\widehat{v}({\boldsymbol{k}})\int_{\mathbb{R}^{d}}v({\boldsymbol{x}})\,\big(k_{1}^{\tilde{s}}+\ldots+k_{d}^{\tilde{s}}\big)H_{\boldsymbol{k}}({\boldsymbol{x}})\,{\rm d}\gamma({\boldsymbol{x}})$
$=2\Gamma(-\tilde{s})\sum_{{\boldsymbol{k}}\in{\mathbb{N}}_{0}^{d}}\big(k_{1}^{\tilde{s}}+\ldots+k_{d}^{\tilde{s}}\big)|\widehat{v}({\boldsymbol{k}})|^{2}.$
@@
$\widehat{v}({\boldsymbol{k}})$ $=\int_{{\mathbb{R}}^{d}}D^{\boldsymbol{\alpha}}f({\boldsymbol{x}})H_{\boldsymbol{k}}({\boldsymbol{x}}){\rm d}\gamma({\boldsymbol{x}})$
@@
$f]_{W_{2}^{s}(\mathbb{R}^{d},\gamma)}^{2}=2\Gamma(-\tilde{s})\sum_{{\boldsymbol{k}}\in{\mathbb{N}}_{0}^{d}}\big(k_{1}^{\tilde{s}}+\ldots+k_{d}^{\tilde{s}}\big)\frac{({\boldsymbol{\alpha}}+{\boldsymbol{k}})!}{{\boldsymbol{k}}!}|\widehat{f}({\boldsymbol{\alpha}}+{\boldsymbol{k}})|^{2}.$
@@ 3 Approximation for Gaussian Sobolev spaces
Let $\Omega$ be a domain in ${\mathbb{R}}^{d}$, and let $Y$ be a Banach space of functions defined on $\Omega$. Let $F\subset Y$ be a compact set and let $n\in\mathbb{N}$. Given sampling points $\{{\boldsymbol{x}}_{i}\}_{i=1}^{n}\subset\Omega$, we aim to approximately recover a function $f\in F$ from the sampled values $\{f({\boldsymbol{x}}_{i})\}_{i=1}^{n}$ by means of a linear sampling algorithm defined by
@@
where $\{\varphi_{i}\}_{i=1}^{n}$ is a family of functions in $Y$. For the sake of generality, we permit repetitions both in the set of sampling points $\{{\boldsymbol{x}}_{i}\}_{i=1}^{n}$ and in the collection of functions $\{\varphi_{i}\}_{i=1}^{n}$. For each $n\in\mathbb{N}$, the sampling $n$-width of the set $F$ in $Y$ is defined as
@@
Let $T$ be a linear bounded operator between two Banach spaces $X$ and $Y$. The $n$th Kolmogorov number of the linear operator $T$ is defined as
$d_{n}(T)=\inf_{L_{n-1}}\sup_{\|x\|_{X}\leq 1}\inf_{y\in L_{n-1}}\|Tx-y\|_{Y}.$
Here the outer infimum is taken over all linear subspaces $L_{n-1}$ of dimension (n1n-1) in $Y$. The $n$th approximation number of the linear operator $T$ is defined as
$a_{n}(T):=\inf\big\{\|T-A\|:\ A\in\mathcal{L}(X,Y),\ \ \text{rank}(A)<n\big\}\,,\qquad n\in{\mathbb{N}}\,.$
The Kolmogorov and approximation numbers belong to the class called $s$-numbers. If $B_{X}$ is the closed unit ball in $X$, then
$d_{n+1}(T)=d_{n}(T(B_{X}),Y),\ \ a_{n+1}(T)=\lambda_{n}(T(B_{X}),Y).$
For a detailed exposition of the theory of widths and $s$-numbers, we refer the reader to the monographs [15, 13, 14].
Denote by $\tilde{L}_{q}({\mathbb{U}}^{d})$ and $\tilde{W}^{\alpha}_{p}({\mathbb{U}}^{d})$ the subspaces of $L_{q}({\mathbb{U}}^{d})$ and $W^{s}_{p}({\mathbb{U}}^{d})$, respectively, of all functions $f$ which can be extended to the whole ${\mathbb{R}}^{d}$ as $\mathbb{U}$}-periodic functions in each variable. The following proposition is crucial for the formulation of our main results.
@@
Choose $q_{*}$ and $p_{*}$ such that $1\leq q<q_{*}<p_{*}<p$. When ${\boldsymbol{x}}\in{\mathbb{I}}^{d}_{\theta,{\boldsymbol{k}}}$ we have $e^{\frac{|{\boldsymbol{x}}|^{2}}{2}}\leq Ce^{\frac{p|{\boldsymbol{k}}|^{2}}{2p_{*}}}$. Therefore,
$\|f_{\boldsymbol{k}}(\cdot+{\boldsymbol{k}})\|_{W_{p}^{\bar{s}}({\mathbb{I}}^{d}_{\theta})}\leq Ce^{\frac{|{\boldsymbol{k}}|^{2}}{2p_{*}}}\|f\|_{W^{s}_{p}({\mathbb{R}}^{d},\gamma)}.$ (3.2)
@@
$\int_{{\mathbb{I}}^{d}_{\theta}}\int_{{\mathbb{I}}^{d}_{\theta}}\frac{\big|D^{\boldsymbol{\alpha}}f_{\boldsymbol{k}}({\boldsymbol{x}}+{\boldsymbol{k}})-D^{\boldsymbol{\alpha}}f_{\boldsymbol{k}}({\boldsymbol{y}}+{\boldsymbol{k}})\big|^{p}}{|{\boldsymbol{x}}-{\boldsymbol{y}}|^{d+\tilde{s}p}}{\rm d}{\boldsymbol{x}}{\rm d}{\boldsymbol{y}}$
$\leq C_{d,t_{0}}\int_{{\mathbb{I}}^{d}_{\theta,{\boldsymbol{k}}}}\int_{{\mathbb{I}}^{d}_{\theta,{\boldsymbol{k}}}}\big|D^{\boldsymbol{\alpha}}f_{\boldsymbol{k}}({\boldsymbol{x}})-D^{\boldsymbol{\alpha}}f_{\boldsymbol{k}}({\boldsymbol{y}})\big|^{p}e^{-\frac{|{\boldsymbol{x}}|^{2}+|{\boldsymbol{y}}|^{2}}{4}\big(1-\frac{1-e^{-t_{0}}}{1+e^{-t_{0}}}\big)}K_{p\tilde{s}}({\boldsymbol{x}},{\boldsymbol{y}}){\rm d}{\boldsymbol{x}}{\rm d}{\boldsymbol{y}}$
$\leq C_{d,t_{0}}\int_{{\mathbb{I}}^{d}_{\theta,{\boldsymbol{k}}}}\int_{{\mathbb{I}}^{d}_{\theta,{\boldsymbol{k}}}}\big|D^{\boldsymbol{\alpha}}f_{\boldsymbol{k}}({\boldsymbol{x}})-D^{\boldsymbol{\alpha}}f_{\boldsymbol{k}}({\boldsymbol{y}})\big|^{p}e^{\frac{|{\boldsymbol{x}}|^{2}+|{\boldsymbol{y}}|^{2}}{4}\big(1+\frac{1-e^{-t_{0}}}{1+e^{-t_{0}}}\big)}K_{p\tilde{s}}({\boldsymbol{x}},{\boldsymbol{y}}){\rm d}\gamma({\boldsymbol{x}}){\rm d}\gamma({\boldsymbol{y}})$
$\leq C_{d,t_{0}}e^{\frac{p|{\boldsymbol{k}}|^{2}}{2p_{*}}\big(1+\frac{1-e^{-t_{0}}}{1+e^{-t_{0}}}\big)}\|f\|_{W^{s}_{p}({\mathbb{R}}^{d},\gamma)}^{p}.$
From this and (3.2) we find that
$\|f_{\boldsymbol{k}}(\cdot+{\boldsymbol{k}})\|_{W^{s}_{p}({\mathbb{I}}^{d}_{\theta})}\leq C_{d,t_{0}}e^{\frac{|{\boldsymbol{k}}|^{2}}{2p_{*}}\big(1+\frac{1-e^{-t_{0}}}{1+e^{-t_{0}}}\big)}\|f\|_{W^{s}_{p}({\mathbb{R}}^{d},\gamma)}.$
@@
$\|f_{{\boldsymbol{k}}}(\cdot+{\boldsymbol{k}})\varphi_{{\boldsymbol{k}}}(\cdot+{\boldsymbol{k}})\|_{W^{s}_{p}({\mathbb{I}}^{d}_{\theta})}$ $\leq C_{t_{0}}\|f_{{\boldsymbol{k}}}(\cdot+{\boldsymbol{k}})\|_{W^{s}_{p}({\mathbb{I}}^{d}_{\theta})}\left\|{\varphi_{\boldsymbol{k}}(\cdot+{\boldsymbol{k}})}\right\|_{C^{\kappa}({\mathbb{R}}^{d})}$
$\leq C_{d,t_{0}}e^{\frac{|{\boldsymbol{k}}|^{2}}{2p_{*}}\big(1+\frac{1-e^{-t_{0}}}{1+e^{-t_{0}}}\big)}\|f\|_{W^{s}_{p}({\mathbb{R}}^{d},\gamma)}.$